# HyperPg - Prototypical Gaussians on the Hypersphere for Interpretable Deep Learning

### v2

@@ 7 Conclusion
This work introduces HyperPg, a new prototype representation learning a probability distribution on the surface of a hypersphere in latent space. HyperPg prototypes adapt to the variance of clusters in latent space and improve training time and accuracy compared to other prototype formulations. HyperPgNet leverages HyperPg to learn human-defined concept prototypes instead of black-box optimized prototypes. The combination of probabilistic prototypes on the hypersphere and concept-aligned prototypes allows HyperPgNet to outperform other prototype learning approaches with regards to accuracy and interpretability. One limitation HyperPgNet faces are slightly higher computational requirements due to the inclusion of concept annotations and $L_{\mathrm{RRC}}$ during training. However, this is offset by faster convergence. Coupled with the increased transparency and interpretability of the model, this makes HyperPgNet a strong contender for scenarios with higher requirements for model trust and safety, like medical applications or human-robot interaction.
Reproducibility Statement
For reproducing the experiments, please refer to the following sections: Sec. 4 describes the required loss functions and the overall network architecture. Subsection 3.3 details the activation function of the newly introduced HyperPg prototypes. Appendix B provides detailed information on the experiment implementations, including training hyperparameters, processing of the datasets and model implementation details. Furthermore, we will publish the code on GitHub after the double-blind review period.
Appendix A Adapting HyperPg to other Probability Distributions
Subsection 3.3 defines HyperPg prototypes ${\bm{p}}^{H}=(\bm{{\alpha}},\mu,\sigma)$ as a Gaussian Distribution with mean $\mu$ and std $\sigma$ of cosine similarities around an anchor vector $\alpha$. This idea of learning a distribution of cosine similarity values around an anchor $\bm{{\alpha}}$ can be adapted to other distributions. This sections introduces some potential candidates. As early experiments on the CUB-200-2011 dataset showed no significant difference in performance, these sections are relegated to the appendix.
A.1 Cauchy Distribution
@@
Fig. 9 illustrates the PDF of the Gaussian and Cauchy distributions with $\mu=x_{0}=1$ and $\sigma=\gamma=0.2$, i.e., the main probability mass is aligned with the anchor $\bm{{\alpha}}$. The Gaussian distributions PDF quickly approaches zero and stays near constant. This could potentially cause vanishing gradient issues during training. The heavier tails of the Cauchy distribution ensure that for virtually the entire value range of the cosine similarity, gradients could be propagated back through the model. However, experiments on CUB-200-2011 showed no significant performance difference between using HyperPg with the Gaussian or Cauchy distribution.
@@ A.2 Truncated Distributions
Fig. 10 shows the test accuracy of three HyperPgNet models with Gaussian, truncated Gaussian and truncated Cauchy distribution on the CUB-200-2011 dataset. While difference in test performance and learning speed were minimal on the CUB-200-2011 dataset, further exploration is necessary, as other experiments showed that the concept-alignment on CUB-200-2011 dominates the learning process, lessening the influence of the prototypes.
Figure 10: CUB-200-2011 Test Accuracy of HyperPgNet with different probability distributions. The difference in performance is only marginal.
A.3 von Mises-Fisher Distribution
The von Mises-Fisher distribution (vMF) is the analogue of the Gaussian distribution on the surface of a hypersphere Hillen et al. (2017). The density function $f_{d}$ of the vMF distribution for a $D$-dimensional unit-length vector ${\bm{v}}$ is defined as
@@
with mean vector $\bm{{\alpha}}$, scalar concentration parameter $\kappa$ and normalization constant $C_{d}(\kappa)$. The normalization constant $C_{d}(\kappa)$ is a complex function and difficult to compute for higher dimensions, which is why, for example, Tensorflow [^1] Tensorflow API Documentation - Accessed 2024-09-20 only supports the vMF distribution for $D\leq 5$. However, the vMF distribution is a viable similarity measure when using the unnormalized density function with $C_{d}(\kappa)=1$. Working with unnormalized densities highlights the relationship between the normal distribution and the vMF distribution.
@@
Eq. 23 also shows the relationship to the HyperPg similarity with an untruncated Gaussian distribution and prototype mean activation $\mu=1$. Fig. 11 presents a simulation of the vMF distribution on a 3D sphere. While both the vMF distribution and HyperPg activation can produce a spherical, gaussian-like activation pattern on the surface of a hypersphere, the vMF distribution cannot produce the ring pattern shown in Fig. 2. The ring pattern produced by adapting HyperPg’s mean similarity $\mu$ could be approximated by a mixture of vMF distributions.
@@
A.4 Fisher-Bingham Distribution
@@
For a $D$ dimensional space, the FB distribution is by a $D\times D$ matrix ${\bm{A}}$ of orthogonal vectors $(\bm{{\alpha}}_{1},\bm{{\alpha}}_{2},\ldots,\bm{{\alpha}}_{D})$, concentration parameter $\kappa$ and ellipticity factors $[\beta]_{2:D}$ where $\sum_{j=2}^{D}\beta_{j}=1$ and $0\leq 2|\beta_{j}|<\kappa$. The FB unnormalzied PDF is defined as
@@
The FB distribution’s main advantage is the elliptic form of the distribution on the surface of the hypersphere, offering higher adaptability than the other formulations (see Fig. 12. However, the parameter count and constraints are higher.
@@ A.5 Mixture Models
Appendix B Implementation Details
B.1 Data Preprocessing
@@
In contrast to prior work (e.g. Chen et al., 2019; Rymarczyk et al., 2020; Ukai et al., 2023) the experiment used an online augmentation process, resulting in 30 training images per class and epoch. The input images were first resized to a resolution of $224\times 224$ without cropping to bounding box annotations. The augmentations consisted of RandomPerspective, RandomHorizontalFlip and RandomAffine.
B.2 Hyperparameters
@@
The prototypical part networks ProtoPNet and HyperPgNet use a convolutional neck after the feature encoder and work on a latent feature map of size $7\times 7\times 256$. The models were trained with a minibatch size of 96 images using AdamW optimizer with learning rate 1e-4 and weight decay 1e-4.
B.3 Compute Resources
@@
All experiments were performed on a workstation with a single NVIDIA RTX 3090 GPU (24 GB VRAM) per model.
B.4 Model Implementation
The models were implement with different feature encoding backbones. ConvNeXt-tiny (Liu et al., 2022) and ResNet50 (He et al., 2016) with pretrained weights on ImageNet provide CNN based backbones. As a transformer based backbone MiT-B4 (Xie et al., 2021) as provided by HuggingFace is used. MiT-B4 is pretrained on ImageNet and finetuned on ADE20k, Cityscapes and COCO-stuff.
The models are implemented as follows:
  • The black box baseline model uses the pretrained feature encoder as backbone followed by the classification head: a single linear output layer and softmax activation.
  • Concept Bottleneck Models (Koh et al., 2020) integrate the concept bottleneck, a linear layer with one neuron per concept, between the backbone and classification head.
  • ProtoPNet (Chen et al., 2019) use a $L_{2}$ based prototype layer between backbone and classification head.
  • ProtoPNet + HyperPg use a HyperPg based prototype module between backbone and classification head.
  • HyperPgNet $-{\mathcal{L}}_{\mathrm{RRC}}$ uses a HyperPg prototype module between backbone and classification head. The prototypes are trained with concept alignment, but without the ${\mathcal{L}}_{\mathrm{RRC}}$ loss.
  • HyperPgNet $+{\mathcal{L}}_{\mathrm{RRC}}$ uses a HyperPg prototype module between backbone and classification head. The prototypes are trained with concept alignment and including the ${\mathcal{L}}_{\mathrm{RRC}}$ loss.
Appendix C Extended Interpretability Analysis
C.1 Latent Space Structure
Prototype learning is based on learning structures in the latent space. HyperPg specifically learns Gaussian distributions on the surface of a hypersphere in a high-dimensional latent space. Dimensionality reduction techniques like UMAP (McInnes et al., 2018) aim to preserve global and local structures from a high dimensional space when projecting into a low dimensional one. UMAP supports multiple distance metrics, including hyperspheric manifold distances, thus retaining some of the high dimensional structure when projecting onto a 3D sphere.
Fig. 13 illustrates UMAP projections of HyperPg concept-aligned prototypes trained on Stanford Cars. This visualization indicates that HyperPgNet is able to disentangle the different concepts in latent space, as the projection shows no overlap of the different clusters. When trained with $L_{\mathrm{RRC}}$, the prototypes are packed closer together. This could indicate, that the chosen hyper parameter of 20 prototypes per concept is higher than required as not all concept embeddings have the same diversity in the latent space.
Fig. 14 illustrates one UMAP projection of HyperPg concept-aligned prototypes trained on CUB-200-2011. In comparison to Stanford Cars, the latent space appears less structured. This could explain the higher difficulty associated with this dataset.
(a) HyperPgNet - $L_{\mathrm{RRC}}$
(b) HyperPgNet + $L_{\mathrm{RRC}}$
Figure 13: UMAP projection of HyperPg concept prototypes learned on Stanford Cars. Each dot represents one prototype with the color indicating a concept.
Figure 14: UMAP Projection of HyperPg concept prototypes learned on CUB-200-2011.

### v3

@@ 7 Conclusion
This work provides comprehensive overview of different prototype formulations, with both probabilistic and point-based measurements, in an Euclidean or hyperspherical latent space. We introduce HyperPG, a new prototype representation based on Gaussian distributions on the surface of a hypersphere. Our experiments showcase the differences in performance and robustness to hyperparameters across different datasets and backbone architectures for these prototype formulations. Our results demonstrate the performance advantage of hypersherical prototypes and probabilistic prototypes in contrast to the commonly used Euclidean formulation, with no loss to the inherent interpretability.
Future work could explore further refinements to the probabilistic prototype formulations, including adaptive mechanisms for prototype selection or the integration of additional probabilistic models such as Mixture Models or Bayesian approaches like MGProto (Wang et al., 2024b) to further enhance interpretability and performance.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Extended Interpretability Analysis
A.1 Potential Problems with Prototype Activation Maps
One common visualization technique originally proposed by ProtoPNet is a a direct overlay of the prototype activation map (PAM). In this manner, the latent similarity map for each prototype is scaled back up to the original input resolution and overlayed over the image. A box is drawn around the 95th quantile to mark the highest activated image region. This technique has some obvious flaws: The latent similarity map is usually very low resolution, e.g., for ResNet50 $7\times 7$, unless a segmentation model such as DeeplabV3 (Chen et al., 2017) is used. Any perceived gradient in the overlay is likely an artifact from the upscaling operation. Additionally, there exist robustness problems with this visualization (Sacha et al., 2024). By modifying parts of the background or adding noise to the image, the activation map can change, most likely due to the spatial pooling in the convolution pyramids.
Appendix B Adapting HyperPG to other Probability Distributions
We define HyperPG prototypes ${\bm{p}}^{H}=(\bm{{\alpha}},\mu,\sigma)$ as a Gaussian Distribution with mean $\mu$ and variance $\sigma^{2}$ of cosine similarities around an anchor vector $\alpha$. This idea of learning a distribution of cosine similarity values around an anchor $\bm{{\alpha}}$ can be adapted to other distributions. This sections introduces some potential candidates. As early experiments on the CUB-200-2011 dataset showed no significant difference in performance, these sections are relegated to the appendix.
B.1 Cauchy Distribution
@@
Figure 18 illustrates the PDF of the Gaussian and Cauchy distributions with $\mu=x_{0}=1$ and $\sigma=\gamma=0.2$, i.e., the main probability mass is aligned with the anchor $\bm{{\alpha}}$. The Gaussian distributions PDF quickly approaches zero and stays near constant. This could potentially cause vanishing gradient issues during training. The heavier tails of the Cauchy distribution ensure that for virtually the entire value range of the cosine similarity, gradients could be propagated back through the model. However, experiments on CUB-200-2011 showed no significant performance difference between using HyperPG with the Gaussian or Cauchy distribution.
@@ B.2 Truncated Distributions
Figure 19 shows the test accuracy of three HyperPGNet models with Gaussian, truncated Gaussian and truncated Cauchy distribution on the CUB-200-2011 dataset. While difference in test performance and learning speed were minimal on the CUB-200-2011 dataset, further exploration is necessary, as other experiments showed that the concept-alignment on CUB-200-2011 dominates the learning process, lessening the influence of the prototypes.
Figure 19: CUB-200-2011 Test Accuracy with HyperPG prototypes using different PDFs.
B.3 von Mises-Fisher Distribution
The von Mises-Fisher distribution (vMF) is the analogue of the Gaussian distribution on the surface of a hypersphere (Hillen et al., 2017). The density function $f_{d}$ of the vMF distribution for a $D$-dimensional unit-length vector ${\bm{v}}$ is defined as
@@
with mean vector $\bm{{\alpha}}$, scalar concentration parameter $\kappa$ and normalization constant $C_{d}(\kappa)$. The normalization constant $C_{d}(\kappa)$ is a complex function and difficult to compute for higher dimensions, which is why, for example, Tensorflow [^1] Tensorflow API Documentation - Accessed 2025-01-10 only supports the vMF distribution for $D\leq 5$.
However, the vMF distribution is a viable similarity measure when using the unnormalized density function with $C_{d}(\kappa)=1$. Working with unnormalized densities highlights the relationship between the normal distribution and the vMF distribution.
@@
Equation 21 also shows the relationship to the HyperPG similarity with an untruncated Gaussian distribution and prototype mean activation $\mu=1$. Figure 20 presents a simulation of the vMF distribution on a 3D sphere. While both the vMF distribution and HyperPG activation can produce a spherical, gaussian-like activation pattern on the surface of a hypersphere, the vMF distribution cannot produce the ring pattern shown in Figure 3. The ring pattern produced by adapting HyperPG’s mean similarity $\mu$ could be approximated by a mixture of vMF distributions.
@@
B.4 Fisher-Bingham Distribution
@@
For a $D$ dimensional space, the FB distribution is by a $D\times D$ matrix ${\bm{A}}$ of orthogonal vectors $(\bm{{\alpha}}_{1},\bm{{\alpha}}_{2},\ldots,\bm{{\alpha}}_{D})$, concentration parameter $\kappa$ and ellipticity factors $[\beta]_{2:D}$ where $\sum_{j=2}^{D}\beta_{j}=1$ and $0\leq 2|\beta_{j}|<\kappa$. The FB’s unnormalized PDF is defined as
@@
The FB distribution’s main advantage is the elliptic form of the distribution on the surface of the hypersphere, offering higher adaptability than the other formulations (see Figure 21. However, the parameter count and constraints are higher.
@@ B.5 Mixture Models
Appendix C Scaled Dot-Product Prototype
This section does not reiterate the full attention mechanism. We recommend to the interested reader to refer to the original work (Vaswani et al., 2017). Rather, we will take a high-level view of the main equation and explain how we apply it to a prototype formulation.
Vaswani et al. (2017) define the scaled dot-product attention for some input matrices query ${\bm{Q}}$, key ${\bm{K}}$ and value ${\bm{V}}$ as
$\mathrm{Attention}({\bm{Q}},{\bm{K}},{\bm{V}})=\mathrm{softmax}\left(\frac{{\bm{Q}}{\bm{K}}^{\top}}{\sqrt{d_{k}}}\right){\bm{V}}$ (28)
with $d_{k}$ being the number of dimensions for the key matrix ${\bm{K}}$.
We can use the scaled dot-product to define a new prototype similarity function
$s_{\mathrm{sdot}}({\bm{z}}|{\bm{p}})=\frac{{\bm{z}}^{\top}{\bm{p}}}{\sqrt{D}}$ (29)
for a latent patch ${\bm{z}}$ and prototype vector ${\bm{p}}$ with dimensionality $D$. In contrast to the cosine similarity, the vectors are not normalized to unit length, which preserves more information. However, dividing by the square root of the number of dimensions $D$ acts as a heuristic regarding vector length. Intuitively we can think about it as instead of a hard projection onto the surface of the unit-radius hypersphere, all points are moved towards the surface. Figure 22 illustrates this scaling. In this manner, the angle between vectors becomes a meaningful measurement, while also retaining some information encoded by the vector length.
Figure 22: Illustration of Cosine Similarity and Scaled Dot-Product
Appendix D Implementation Details
D.1 Data Preprocessing
@@
In contrast to prior work (e.g. Chen et al., 2019; Rymarczyk et al., 2020; Ukai et al., 2023) the experiment used an online augmentation process, resulting in 30 training images per class and epoch. The input images were cropped to the bounding box annotations, then resized to a resolution of $224\times 224$. The augmentations consisted of RandomPerspective, RandomHorizontalFlip and ColorJitter.
D.2 Hyperparameters
@@
The prototypical part networks ProtoPNet and HyperPGNet use a convolutional neck after the feature encoder and work on a latent feature map of size $7\times 7\times 128$. The models were trained with a minibatch size of 48 images using AdamW optimizer with learning rate 1e-4 and weight decay 1e-4.
D.3 Compute Resources
@@
All experiments were performed on a desktop workstation with a single NVIDIA RTX 3070Ti GPU (8 GB VRAM) per model. On CUB-200-2011 the training duration was roughly 40 seconds per epoch, regardless of the used prototype formulation.
D.4 Model Implementation
The experiments used various feature encoding backbones such as ResNet50 (He et al., 2016), DenseNet121 (Huang et al., 2017) and ViT-B16 (Dosovitskiy, 2020) with pretrained weights on ImageNet. All models use a single linear output layer as a classification head.
The prototypical networks follow the architecture proposed by ProtoPNet (Chen et al., 2019): After the pretrained feature encoder, a bottleneck consisting of a Convolution Layer, ReLU activation, Convolution Layer and Sigmoid activation reduces the dimensions of the feature map to the prototype space. For example, the ResNet50 encoder produces a feature map of $H\times W\times D$ shape $7\times 7\times 2048$. The bottleneck produces a dimension reduction to shape $7\times 7\times 128$.
The prototype modules produce the prototype similarity scores, which are passed through a global max pooling, before being passed to the classification head.
Appendix E Parameter Distribution of HyperPG
To evaluate, if HyperPG actually uses the probabilstic components, we look at the learned mean $\mu$ and standard deviation $\sigma$ values after training on CUB200-2011.
Figure 23 plots the mean and sigma values. The red star marks the theoretical solution if it would learn something similar to the pure Cosine prototypes. The pattern demonstrates, that HyperPG actually learns different parameters for the Gaussian components of each prototype. Interestingly, two modes at $\sigma=0.5$ and $\sigma=0.4$ appear with a distinct line shape, which warrant future exploration.
Figure 23: HyperPG learned mean and sigma values.